You and your friend decide to get your cars inspected. You are informed that 85​% of cars pass inspection. If the event of your​ car's passing is independent of your​ friend's car, ​a) What is the probability that your car passes​ inspection? ​b) What is the probability that your car​ doesn't pass​ inspection? ​c) What is the probability that both of the cars​ pass? ​d) What is the probability that at least one of the two cars​ passes?
Answer #1

a) the probability that your car passes inspection is 0.85

b) the probability that your car doesn’t pass inspection

is (1-0.85) = 0.15

c) the probability that both of the cars pass is

(0.85*0.85) = 0.7225

d) the probability that at lease on of the two cars passes is

0.85 + 0.85 - (0.85 *0.85) = 0.9775